OpenStudy (anonymous):

The point (3, -4) is on the terminal side of an angle θ . What is cos θ?

OpenStudy (misty1212):



OpenStudy (misty1212):

find the hypotenuse of that right triangle, lets call it \(h\) it will be \(\cos(\theta)=\frac{3}{h}\)

OpenStudy (misty1212):

the hypotenuse you can do with your eyeballs if you remember the mother of all right triangles, the 3 - 4 - 5 right triangle or you could use pythagoras if you do not remember that one

OpenStudy (anonymous):

Through high school ive had sucky math teachers, so i barely learned anything. So, im having to take online classes and im a visual learner and dont know how to do any of this. Its so confusing

OpenStudy (misty1212):

what you need to know for this is that the cosine is the first coordinate divided by the length of the hypotenuse

OpenStudy (misty1212):

you are told that the first coordinate is 3 so that is your numerator also the other leg of the triangle has length 4 that means, by pythangoras, since \[3^2+4^2=h^2\] you have \[9+16=h^2\\ 25=h^2\\ 5=h\]

OpenStudy (anonymous):

here are the answer choices a) - 3/5 b) 3/5 c) -3/4 D) 3/4 E) 3

OpenStudy (misty1212):

ok which one is it?

OpenStudy (anonymous):

b

OpenStudy (misty1212):

yes